OpenStudy (anonymous):

log3 + log√3 9x = -2

OpenStudy (raden):

hint : log√3 9x = log(3)^1/2 9x = log3 (9x)^2 = log3 81x^2

OpenStudy (raden):

can u continue it

OpenStudy (anonymous):

i got 1/81. it supposed to be 1/9. can u show me the calculation?

OpenStudy (raden):

wait, is the log using 10 as the base ?

OpenStudy (raden):

i think yes, so miss my the first hint (that was wrong)

OpenStudy (anonymous):

okay. my bad,im sorry . the ques should be log (base=3) x + log (base=√3) 9x = -2

OpenStudy (raden):

ah, still using my hint above :)

OpenStudy (raden):

ok, now we have equation : log3 x + log3 81x^2 = -2 simplify again, becomes log3 (81x^2 * x) = -2 log3 (81x^3 ) = -2 convert to exponent form : 81x^3 = 3^-2 or 81x^3 = 1/9 x^3 = 1/(9 * 81) x^3 = 1/729 x^3 = (1/9)^3 thus, the value of x = 1/9